2. You performed an X-ray spectroscopy experiment and observed the emission from unknown element A to be 21×1016 Hz and the emission from unknown element B to be 71.5×1016 Hz. It is known that X-ray frequency increases with atomic number and fits the quadratic function of Y=X2. If Unknown A is Neon, what is the likely identity of element B?


1
Expert's answer
2021-10-29T02:30:30-0400

Unknown B is likely Argon


Explanation:-

"\frac{Element B}{Element A}" ="\frac{71.5}{21}=3.4"


But the quadratic function Y=X2

Y=3.4×2=6.8

Frequency =6.8 Hence the frequency of Argon